OpenStudy (anonymous):

\[ \sum_{n=1}^6 {n^2}+\sum_{n=1}^6 \frac{n(n+1)}{2} = \cdot\cdot\cdot \]

OpenStudy (helder_edwin):

\[ \large =\sum_{n=1}^6n^2+\frac{1}{2}\sum_{n=1}^6(n^2+n)= \frac{3}{2}\sum_{n=1}^6n^2+\frac{1}{2}\sum_{n=1}^6n= \]